OpenStudy (anonymous):

A blueprint for a house has a scale of 1 : 30. A wall in the blueprint is 7 in. What is the length of the actual wall?

OpenStudy (anonymous):

7 x 30

OpenStudy (luffingsails):

A scale of 1:30 in this case says: "for every 1 inch on the blueprint it is 30 inches in real life." So if I measure 7 inches on the blueprint, I can use the formula: 1/30 = 7/x and then solve for x. x = 30 * 7 = 210 inches.